# NV-Embed: Improved Techniques for Training LLMs as Generalist Embedding Models

### v1

@@ Appendix A Implementation Details
Table 6: Instructions and number of samples used for each training dataset. For an apples-to-apples comparison, we use the same instructions as in Wang et al. (2023b) for shared training datasets.
@@
ArguAna Given a claim, find documents that refute the claim 16k
@@
100k, 200k
SQUAD Given a question, retrieve Wikipedia passages that answer the question 100k
StackExchange Given a question paragraph at StackExchange, retrieve a question duplicated paragraph 80k
Natural Question Given a question, retrieve Wikipedia passages that answer the question 100k
HotpotQA Given a multi-hop question, retrieve documents that can help answer the question 50k
@@
FiQA2018 Given a financial question, retrieve user replies that best answer the question 5k
BioASQ Given a question, retrieve detailed question descriptions that are duplicates to the given question 2k
STS12, STS22, STSBenchmark Retrieve semantically similar text. 40k
AmazonCounterfactualClassification Classify a given Amazon customer review text as either counterfactual or not-counterfactual 10k
AmazonReviewsClassification Classify the given Amazon review into its appropriate rating category 20k
Banking77Classification Given a online banking query, find the corresponding intents 10k
EmotionClassification Classify the emotion expressed in the given Twitter message into one of the six emotions:anger, 16k
@@
ImdbClassification Classify the sentiment expressed in the given movie review text from the IMDB dataset 15k
MTOPIntentClassification Classify the intent of the given utterance in task-oriented conversation 10k
ToxicConversationsClassification Classify the given comments as either toxic or not toxic 40k
TweetSentimentExtractionClassification Classify the sentiment of a given tweet as either positive, negative, or neutral 40k
ArxivClusteringP2P Identify the main and secondary category of Arxiv papers based on the titles and abstracts 25k
ArxivClusteringS2S Identify the main and secondary category of Arxiv papers based on the titles 25k
BiorxivClusteringP2P Identify the main category of Biorxiv papers based on the titles and abstracts 25k
BiorxivClusteringS2S Identify the main category of Biorxiv papers based on the titles 15k
MedrxivClusteringP2P Identify the main category of Medrxiv papers based on the titles and abstracts 15k
MedrxivClusteringS2S Identify the main category of Medrxiv papers based on the titles 15k
TwentyNewsgroupsClustering Identify the topic or theme of the given news articles 10k
Table 7: Instructions used for evaluation on the MTEB benchmark. “STS*” indicates we use the same instructions for all the STS tasks. For an apples-to-apples comparison, we use the same instructions as in Wang et al. (2023b).
@@
ArguAna Given a claim, find documents that refute the claim
@@
NFCorpus Given a question, retrieve relevant documents that best answer the question
Natural Question Given a question, retrieve Wikipedia passages that answer the question
@@
Touche2020 Given a question, retrieve detailed and persuasive arguments that answer the question
TREC-COVID Given a query, retrieve documents that answer the query

### v2

@@ Appendix B Experimental Details and Instruction Templates for Training and Evaluation
Table 9: Instructions and number of samples used for each training dataset.
@@
ArguAna Given a claim, retrieve documents that support or refute the claim 16k
@@
500k, 500k
SQUAD Given a question, retrieve passages that answer the question 87k
StackExchange Given a web search query, retrieve relevant passages that answer the query 80k
Natural Question Given a question, retrieve passages that answer the question 100k
HotpotQA Given a multi-hop question, retrieve documents that can help answer the question 170k
@@
FiQA2018 Given a financial question, retrieve relevant passages that answer the query 5k
BioASQ Given a query, retrieve documents that can help answer the question 2.4k
HoVer Given a claim, retrieve documents that support or refute the claim 17k
Nfcorpus Given a question, retrieve relevant documents that answer the question 3.6k
MIRACL Given a question, retrieve passages that answer the question 2k
Mr.TyDi Given a question, retrieve passages that answer the question 2k
SciFact Given a scientific claim, retrieve documents that support or refute the claim 0.9k
STS12, STS22, STSBenchmark Retrieve semantically similar text. 1.8k, 0.3k, 2.7k
AmazonCounterfactual-Classification Classify a given Amazon customer review text as either counterfactual or not-counterfactual 6k
AmazonPolarity-Classification Classify Amazon reviews into positive or negative sentiment 20k
AmazonReviews-Classification Classify the given Amazon review into its appropriate rating category 40k
Banking77-Classification Given a online banking query, find the corresponding intents 10k
Emotion-Classification Classify the emotion expressed in the given Twitter message into one of the six emotions:anger, 16k
@@
Imdb-Classification Classify the sentiment expressed in the given movie review text from the IMDB dataset 24k
MTOPIntent-Classification Classify the intent of the given utterance in task-oriented conversation 15k
MTOPDomain-Classification Classify the intent domain of the given utterance in task-oriented conversation 15k
MassiveIntent-Classification Given a user utterance as query, find the user intents 11k
MassiveScenario-Classification Given a user utterance as query, find the user scenarios 11k
ToxicConversationsClassification Classify the given comments as either toxic or not toxic 50k
TweetSentimentExtractionClassification Classify the sentiment of a given tweet as either positive, negative, or neutral 27k
Arxiv-Clustering-P2P Identify the main and secondary category of Arxiv papers based on the titles and abstracts 50k
Arxiv-Clustering-S2S Identify the main and secondary category of Arxiv papers based on the titles 50k
Biorxiv-Clustering-P2P Identify the main category of Biorxiv papers based on the titles and abstracts 15k
Biorxiv-Clustering-S2S Identify the main category of Biorxiv papers based on the titles 15k
Medrxiv-Clustering-P2P Identify the main category of Medrxiv papers based on the titles and abstracts 2.3k
Medrxiv-Clustering-S2S Identify the main category of Medrxiv papers based on the titles 2.3k
Reddit-Clustering Identify the main category of Medrxiv papers based on the titles and abstracts 50k
Reddit-Clustering-S2S Identify the main category of Medrxiv papers based on the titles and abstracts 40k
Stackexchange-Clustering Identify the main category of Medrxiv papers based on the titles and abstracts 50k
Stackexchange-Clustering-S2S Identify the main category of Medrxiv papers based on the titles and abstracts 40k
TwentyNewsgroups-Clustering Identify the topic or theme of the given news articles 1.7k
Table 10: Instructions used for evaluation on the MTEB benchmark. “STS*” indicates we use the same instructions for all the STS tasks.
@@
ArguAna Given a claim, retrieve documents that support or refute the claim
@@
NFCorpus Given a question, retrieve relevant documents that answer the question
Natural Question Given a question, retrieve passages that answer the question
@@
Touche2020 Given a question, retrieve passages that answer the question
TREC-COVID Given a query on COVID-19, retrieve documents that answer the query
@@
Appendix C Latent-Attention Visualization
Table 11: Full BEIR and MTEB benchmark
Bge-multilin
gual-gemma2
Gte-Qwen2-
7B-instruct
SFR-Embe
dding-2R
Stella-en-
1.5B-v5
bge-en-icl
NV-Embed-v1 NV-Embed-v2
ArguAna 77.37 64.27 62.34 65.27 82.76 68.21 70.07
ClimateFEVER 39.37 45.88 34.43 46.11 45.35 34.72 45.39
CQADupStack 47.94 46.43 46.11 47.75 47.23 50.51 50.24
DBPEDIA 51.37 52.42 51.21 52.28 50.42 48.29 53.50
FEVER 90.38 95.11 92.16 94.83 91.96 87.77 93.75
FiQA2018 60.04 62.03 61.77 60.48 58.77 63.1 65.73
HotpotQA 83.26 73.08 81.36 76.67 84.98 79.92 85.48
MSMARCO 45.71 45.98 42.18 45.22 46.72 46.49 45.63
NFCorpus 38.11 40.6 41.34 42 40.69 38.04 45.17
Natural 71.45 67 73.96 71.8 73.85 71.22 73.57
QuoraRetrieval 90.04 90.09 89.58 90.03 91.02 89.21 89.04
SCIDOCS 26.93 28.91 24.87 26.64 25.25 20.19 21.90
SciFact 72.05 79.06 85.91 80.09 78.33 78.43 80.13
Touche2020 30.26 30.57 28.18 29.94 29.67 28.38 88.44
TREC-COVID 64.27 82.26 87.28 85.98 78.11 85.88 31.78
BIOSSES 85.74 81.37 87.6 83.11 86.35 85.59 87.42
SICK-R 82.66 79.28 77.01 82.89 83.87 82.8 82.15
STS12 77.71 79.55 75.67 80.09 77.73 76.22 77.89
STS13 87.45 88.83 82.4 89.68 85.98 86.3 88.30
STS14 83.48 83.87 79.93 85.07 82.34 82.09 84.30
STS15 87.63 88.54 85.82 89.39 87.35 87.24 89.04
STS16 86.7 86.49 84.5 87.15 86.54 84.77 86.77
STS17 91.18 88.73 88.93 91.35 91.25 87.42 90.67
STS22 69.02 66.88 67.1 68.1 68.08 69.85 68.12
STSBenchmark 87.25 86.85 83.6 88.23 87.92 86.14 88.41
SummEval 31.2 31.35 30.71 31.49 30.75 31.2 30.70
SprintDuplicateQuestions 90.94 92.82 97.62 96.04 95.06 95.94 97.02
TwitterSemEval2015 79.64 77.96 78.57 80.58 78.54 78.73 81.11
TwitterURLCorpus 86.95 86.59 88.03 87.58 87.19 86.05 87.87
AmazonCounterfactual 89.48 91.31 92.72 92.87 92.88 95.12 94.28
AmazonPolarity 96.9 97.5 97.31 97.16 96.86 97.14 97.74
AmazonReviews 61.6 62.56 61.04 59.36 61.28 55.47 63.96
Banking77 92.53 87.57 90.02 89.79 91.42 90.34 92.42
Emotion 92.97 79.45 93.37 84.29 93.31 91.71 93.38
Imdb 96.66 96.75 96.8 96.66 96.91 97.06 97.14
MassiveIntent 82.05 85.41 85.97 85.83 82.26 80.07 86.10
MassiveScenario 84.4 89.77 90.61 90.2 83.92 81.74 92.17
MTOPDomain 98.61 99.04 98.58 99.01 97.99 96.51 99.25
MTOPIntent 95.51 91.88 91.3 92.78 93.56 89.77 94.37
ToxicConversations 87.34 85.12 91.14 88.76 93.16 92.6 92.74
TweetSentimentExtraction 78.86 72.58 79.7 74.84 79.9 80.6 80.87
Arxiv-P2P 54.91 54.46 54.02 55.44 54.42 53.76 55.80
Arxiv-S2S 50.28 51.74 48.82 50.66 49.17 49.59 51.26
Biorxiv-P2P 52.64 50.09 50.76 50.68 52.32 48.15 54.09
Biorxiv-S2S 49.2 46.65 46.57 46.87 48.38 44.74 49.60
Medrxiv-P2P 45.81 46.23 46.66 46.87 46.13 39.24 46.09
Medrxiv-S2S 44.11 44.13 44.18 44.65 44.2 36.98 44.86
Reddit 56.03 73.55 62.92 72.86 71.2 63.2 71.10
Reddit-P2P 65.83 74.13 72.74 75.27 72.17 68.01 74.94
StackExchange 66.21 79.86 76.48 80.29 81.29 74.99 82.10
StackExchange-P2P 45.74 49.41 48.29 49.57 45.53 42.04 48.36
TwentyNewsgroups 70.44 53.91 66.42 61.43 68.51 60.13 64.82
AskUbuntuDupQuestions 64.59 67.58 66.71 67.33 64.8 67.5 67.46
MindSmallRerank 31.79 33.36 31.26 33.05 30.6 30.82 31.76
SciDocsRR 87.6 89.09 87.29 89.2 86.9 87.26 87.59
StackOverflowDupQuestions 54.9 55.66 55.32 55.25 56.32 56.58 55.79
MTEB Average (56) 69.88 70.24 70.31 71.19 71.24 69.32 72.31
Table 12: Prompt template for short-long matching subgroup.
Brainstorm a list of potentially useful text retrieval tasks.
Here are a few examples for your reference:
- Given a claim about climate change, retrieve documents that support or refute the claim
- Given a job title, search for job descriptions that provide information about the role
Please adhere to the following guidelines:
- Specify the type of query and the type of desired texts.
- Each retrieval task should cover a wide range of queries, and should not be too specific.
- Cover a wide range of query types and desired text types.
Your output must always be a JSON list of strings only, with about 40 elements, and each element corresponds to
a distinct retrieval task in one sentence. Do not explain yourself or output anything else. Be creative!
You have been assigned a retrieval task: {task}
Your mission is to write one text retrieval example for this task in JSON format. The JSON object must
contain the following keys:
- "user_query": a string, a random example of what is provided as specified by the task description.
- "positive_document": a string, a relevant document for the user query.
- "hard_negative_document1": a string, a hard negative document that is irrelevant but appears relevant to the query.
- "hard_negative_document2": a string, another hard negative document that is irrelevant but appears relevant to the query.
Please adhere to the following guidelines:
- The "user_query" should be {query_type}, {query_length}, {clarity}, and diverse in topic. The "user_query" should
not restate the task and just contain what the task description says is provided.
- All documents must be created independent of the query. Avoid copying the query verbatim. It’s acceptable if
some parts of the "positive_document" are not topically related to the query.
- All documents should be at least {num_words} words long.
- The "hard_negative_document1" may contain little useful information, but it should be less useful or
comprehensive compared to the "positive_document".
- The "hard_negative_document2" may should be about a related but different topic.
- Do not provide any explanation in any document on why it is relevant or not relevant to the query.
- Both the query and documents require {difficulty} level education to understand.
Your output must always be a JSON object only, do not explain yourself or output anything else. Be creative!"""
Placeholders:
[…]
Table 13: Prompt template for long-short matching subgroup.
Brainstorm a list of potentially useful text classification tasks.
Please adhere to the following guidelines:
- Tasks should cover a diverse range of domains and task types.
Your output must always be a JSON list of strings only, with about 40 elements, and each element corresponds
to a distinct text classification task in one sentence. Do not explain yourself or output anything else. Be creative!
You have been assigned a text classification task: {task}
Your mission is to write one text classification example for this task in JSON format. The JSON object must
contain the following keys:
- "input_text": a string, the input text specified by the classification task.
- "label": a string, the correct
label of the input text.
- "misleading_label": a string, an incorrect label that is related to the task.
Please adhere to the following guidelines:
- The "input_text" should be {num_words} words and diverse in expression.
- The "misleading_label" must be a valid label for the given task, but not as appropriate as the "label" for the
"input_text".
- Avoid including the values of the "label" and "misleading_label" fields in the "input_text", that would make
the task too easy.
- The "input_text" is {clarity} and requires {difficulty} level education to comprehend.
Your output must always be a JSON object only, do not explain yourself or output anything else. Be creative!
Placeholders:
[…]
Figure 2: Attention over 4096 latents across 8 heads (columns) are visualized for 10 positive and 10 negative reviews (rows) from the AmazonReviewsClassification dataset. The attention weights are mean pooled across tokens. The attention weights reveal that the latents specialize in learning features of queries. The latent indicated by the arrows specialized in learning the positivity of reviews. It has high attention across the positive reviews and low attention across the negative reviews.